OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (campbell_st):

rewrite the equations in point slope form y = 1/2 x -2 y = -9/2 x - 12 use a table of values for each line  graph the lines and find the point of intersection. Substitute the point into each like to check it is a solution. the point is the where the lines have a common x and y value.